OpenStudy (amonoconnor):

Is the following true: *I will post the mathematical expression as the first comment. Any and all help is greatly appreciated!

OpenStudy (amonoconnor):

My question: \[\int\limits_{-\pi/4}^{\pi/4}(x^3+x^4\tan(x))dx = \int\limits_{-\pi/4}^{\pi/4}(u*\tan(u^{1/4})*du/4) \] if u = x^4. ???

OpenStudy (misty1212):

HI!!

OpenStudy (amonoconnor):

Hey misty:)

OpenStudy (misty1212):

i wouldn't think so how did the sum become a product?

OpenStudy (amonoconnor):

Damn.. I thought that might be it. So, would I distribute the dx, and then solve for dx in the du=4x^3*dx part?

OpenStudy (misty1212):

btw the answer is pretty clear without finding any anti derivative, which is almost impossible

OpenStudy (amonoconnor):

I thought about factoring out an x^3... is that it? :/

OpenStudy (misty1212):

i think the point of this question is something else entirely

OpenStudy (misty1212):

look carefully notice anything about the limits of integration?

OpenStudy (misty1212):

let me put in another way you are not going to find a closed form for the anti derivative, ie. you are not going to find some function whose derivative is the integrand it is testing something else entirely

OpenStudy (amonoconnor):

it's a half-pi interval, centered on 0? :/ I'll admit, I'm lost right now...

OpenStudy (misty1212):

want a hint?

OpenStudy (amonoconnor):

:( Please!

OpenStudy (misty1212):

the "centered at zero " part is good

OpenStudy (misty1212):

what kind of function is this? hint, compute \[f(\frac{\pi}{4})\] and \[f(-\frac{\pi}{4})\]

OpenStudy (misty1212):

ok i will stop torturing you the function is odd

OpenStudy (amonoconnor):

OMG!! My professor mentioned that about 2 weeks ago.. vaguely. I see where you're going now!! What's the definition of an odd function?

OpenStudy (misty1212):

\[f(-x)=-f(x)\]

OpenStudy (amonoconnor):

Ah.... so it's just sort of an "understand how these work, and look for the hints when they show up" situation?

OpenStudy (misty1212):

which is clear because \(x^3\)is odd and \(tan(x)\) is odd and \(x^4\) is even, so \[(-x)^3+(-x)^4\tan(-x)=-x^2-x^4\tan(x)=-(x^3+x^4\tan(x))=-f(x)\]

OpenStudy (misty1212):

you integrate an odd function \[\int_{-a}^af(x)dx\]you get \(0\) because of the symmetry

OpenStudy (amonoconnor):

Goddamn you're blowing my mind, and I love it. Gotta... Like, it's reflected around the x-axis because it's Tangent, and since it's from |x| on either side of the y-axis, the two parts are equal to each other, just one's negative? :D

OpenStudy (amonoconnor):

So the answer is 0? :D

OpenStudy (misty1212):

yes, definitely zero think about what it means for a function to be odd

OpenStudy (amonoconnor):

Thank you so much for your help:)

OpenStudy (misty1212):



OpenStudy (amonoconnor):

It's possible to get that algebraically though, with u-sub right?

OpenStudy (misty1212):

area above equals area below actually easiest to see with simplest odd function \(f(x)=x\)

OpenStudy (misty1212):

no

OpenStudy (misty1212):

forget it

OpenStudy (amonoconnor):

... what do I put as "work", just a little not that it's an odd function?

OpenStudy (misty1212):

look, just because you can write down a function doesn't mean you can find its anti derivative it is almost never the case that you can

OpenStudy (misty1212):

all that calc 2 stuff you are doing is just showing off mostly the anti derivative does not have a closed form, the "work" is to say "function is odd, interval is from \(-a\) to \(a\), answer is 0"

OpenStudy (amonoconnor):

Bueno, sounds good to me:)

OpenStudy (misty1212):

that was the entire point of this question, to see if you would realize that

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]

OpenStudy (amonoconnor):

♡ :)